OpenStudy (anonymous):

Yura's trip to Monkville was a distance of 738 miles. If she drove for 11.2 hours, what mph did she average on her trip? Use the appropriate formula and round your answer to the nearest whole number. A. 67 mph B. 66 mph C. 65 mph Yura's trip to Monkville was a distance of 738 miles. If she drove for 11.2 hours, what mph did she average on her trip? Use the appropriate formula and round your answer to the nearest whole number. A. 67 mph B. 66 mph C. 65 mph @Mathematics

OpenStudy (saifoo.khan):

\[speed = \frac{distance}{time}\]

OpenStudy (saifoo.khan):

btw, nice name.

OpenStudy (anonymous):

\[\large speed = \frac{distance}{time}\]

OpenStudy (anonymous):

\[\frac{738}{11.2}\]

OpenStudy (anonymous):

\[\approx 65.89\] so round up to 66mph

OpenStudy (anonymous):

lol thanks :P